A woman was 32years when her son was born. If the sum of their ages is 56years, how old is the woman?
Answer by macston(5194)

S=son's age;W=woman's age
W=S+32 years
S+W=56 years Substitute for W.
S+S+32 years=56 years Subtract 32 years from each side.
2S=24 years Divide each side by 2.
S=12 years ANSWER 1: The son is 12 years old
W=S+32years=12years+32 years=44 years ANSWER 2: The woman is 44 years old.
CHECK:
S+W=56 years
12 years + 44 years=56 years
56 years=56 years